the perimeter of a rectangle is 84 meters. The length is 6 meters longer than twice the width. find he dimensions.
Answer by sofiyac(983)

let the width be x, then the length is 2x+6
perimeter is all sides added together so
x+x+2x+6+2x+6=84 solve for x
6x+12=84 subtract 12 from each side of the equal sign
6x=72 divide each side by 6
x=12 so width is 12 now plug that back into what we said length was in the first sentence, so length=2x+6=2*12+6=24+6=30
width 12, length 30